# AI/ML based Joint Source and Channel Coding for HARQ-ACK Payload

### v1

@@ II System Model
Consider a channel code mapping $k$ information bits to $n$ coded symbols, henceforth identified by the tuple $(k,n)$. An information bit value of 0 corresponds to NACK, while 1 corresponds to ACK. A message sequence $\mathbf{b}_{m}=\{b_{0},b_{1},\ldots b_{k-1}\}\in\{0,1\}^{k}$ consisting of $k$ information bits is jointly encoded and modulated to $n$ real values (or equivalently $n/2$ complex values). The message sequence index is denoted by $m$ (for e.g. if $k=4$, then the message sequence 0101 corresponds to index 5). The complex valued encoded sequence is denoted by $\mathbf{c}_{m}$. Each codeword $\mathbf{c}_{m}$ has norm $\|\mathbf{c}_{m}\|^{2}_{2}=n/2$ and is associated with a prior probability $\pi_{m}$ such that $\sum_{m\in[2^{k}]}\pi_{m}=1$. If the encoding is binary, the intermediate encoded bit sequence is denoted by $\mathbf{e}(\mathbf{b}_{m})$. Denote the codebook by $\mathcal{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots\mathbf{c}_{2^{k}-1}\}$. Subsequently, each codeword $\mathbf{c}_{m}$ is multiplied by a scalar per-codeword power shaping factor $\sqrt{\alpha_{m}}$ (to be introduced in Section IV). In current NR specifications [3gpp.38.212], all codewords are transmitted with equal power i.e. $\alpha_{m}=1$.
@@ II-A AWGN setup
where $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{n/2})$ and $\mathrm{SNR}=g^{2}/\sigma^{2}$.
@@ II-B NR-compliant fading channel setup
Consider the end-to-end (E2E) frequency domain (FD) pipeline for uplink transmission of PUCCH [^3] Note that a similar formulation can be derived for HARQ on PUSCH as well. We focus only on PUCCH for simplicity of the SIMO setup. HARQ-ACK as depicted in Fig. 1. We assume a SIMO setup with $N_{r}$ receive antennas, since PUCCH is transmitted from a single antenna port [johnson20195g]. Under the assumption of coherent reception i.e. excluding PUCCH Format 0, we would also transmit $N_{p}$ Demodulation Reference Symbols (DMRS) resource elements (REs) which we set to 1 w.l.o.g. given a single transmit port. DMRS REs are utilized by the receiver to perform channel estimation (ChEST). Suppose that the DMRS REs are also multiplied by a power shaping factor $\sqrt{\beta_{m}}$. In this paper, we will be considering two cases - no power shaping on the DMRS i.e. $\beta_{m}=1$ and data dependent DMRS power shaping i.e. $\beta_{m}=\alpha_{m}$.
@@
$\mathbf{y}_{l}=\epsilon(g\mathbf{h}_{l}\odot\mathbf{c}_{p+d,m}+\mathbf{n}_{l}),$ (3)
where $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$ and $\epsilon$ captures the overall scaling introduced in receiver processing. For simplicity, we assume that the same $\epsilon$ is applied to all receive antennas.
Denote the received signal (across all receive antennas) by $\mathbf{Y}=[\mathrm{y}_{1},\ldots\mathrm{y}_{N_{r}}]$. Depending on whether we choose to perform coherent or noncoherent decoding, the received signal is passed to a channel estimator module before being input to a decoder. Furthermore, we can define the long term SNR (per RE and per receive antenna) and noise variance (NV) as
$\mathrm{SNR}$ $=\frac{\epsilon^{2}g^{2}}{\epsilon^{2}\sigma^{2}}=\frac{g^{2}}{\sigma^{2}}$ (4)
$\mathrm{NV}$ $=\epsilon^{2}\sigma^{2}.$
@@ III-A Transformer (TF) Architecture
We utilize a transformer (TF) encoder [vaswani2017attention] to encode the input bits. The architecture is illustrated in Fig. 2. The input bit sequence $\mathbf{b}_{m}\in\{0,1\}^{k}$ is divided into $A$ tokens of size $\lceil k/A\rceil$. If $k<A\lceil k/A\rceil$, the first $k$ positions are filled with input bits and the remaining $A\lceil k/A\rceil-k$ positions are set to 0. Subsequently, we BPSK modulate the input bits $\mathbf{m}=(-1)^{\mathbf{b}_{m}}$ before inputting the tokens to the transformer. The transformer encoder consists of $L=6$ layers. Each layer consists of the standard self attention and feedforward networks with residual connections. We utilize an embedding length [^4] Refer https://pytorch.org/docs/stable/generated/torch.nn.TransformerEncoder-Layer.html for meaning of each hyperparameter of $\mathrm{d}_{\mathrm{model}}=128$, feedforward dimension $\mathrm{dim}_{\mathrm{feedforward}}=512$ and $\mathrm{nhead}=8$. The tokens input to the transformer and output by it are passed through Layer Norm [xiong2020layer] normalization. Positional embedding is not added to the input tokens. The transformer outputs $A$ tokens. Each output token is passed through a linear layer $H$ to obtain $\lceil n/A\rceil$ coded symbols per token. The coded symbols computed from the output tokens are concatenated to obtain the codeword of $n$ coded symbols, which can then be normalized to ensure the total codeword power of each codeword is $n$. Finally, the transformer output can be reshaped to obtain $\mathbf{c}_{m}$, an $n/2$ length complex codeword.
@@ III-B Codebook Properties
Since the encoding operations involve computing self-attention on the message tokens, the codebook is non-linear. Non-linearity implies that each codeword will have a different Block Error Rate (BLER)-SNR curve, leading to a per-codeword BLER spread. In order to reduce the per-codeword BLER spread, the message sequence to codeword mapping can be scrambled, such that every message sequence sees the same BLER curve on average.
Since we do not utilize positional embedding, the codebook learnt is group permutation symmetric with a group size of $\lceil k/A\rceil$. This implies we do not need to store the embedding of any group permutations of the input message sequence. Given an input sequence $\mathbf{x}$, we do not need to invoke the transformer to compute the encoding of any sequence that can be written as $\mathbf{Px}$, where $\mathbf{P}\in\{0,1\}^{A\times A}$ and $\mathbf{x}=[\mathbf{m}_{0}^{T}\mathbf{m}_{1}^{T}\ldots\mathbf{m}_{A-1}^{T}]\in\{-1,1\}^{A\times\lceil k/A\rceil}$. This implies that for any input sequence $\mathbf{x}$, there are at most A!1A!-1 orderings of its input tokens [^5] If some of the tokens are identical, then the number of distinct orderings will be smaller. whose encoding can be computed by group permutation of $H(\mathbf{x})$. This property can be exploited to reduce the size of look-up table (LUT) used to store the codebook.
Finally, increasing the number of tokens $A$ (or equivalently reducing the token size $\lceil k/A\rceil$) increases the degree of group permutation symmetry in the codebook. In other words, larger symmetry reduces the size of the LUT required to store the codebook. Increased symmetry also reduces the per-codeword BLER spread arising from codebook non-linearity.
@@ IV-A Entropy PS
Figure 3: Entropy and Learnt PS as a function of $\pi_{m}$
Observe that codewords with high probability of occurrence $\pi_{m}$ have low power $\alpha_{m}$ and vice-versa. Moreover, codewords with the same probability of occurrence have the same codeword power. Since we plot as a function of $\pi_{m}$, Entropy PS is a single dot but the Learnt PS has a spread (if more than one codeword has the same prior probability). Since codewords with low $\pi_{m}$ are seen less frequently in training, this is to be expected. However, for all codewords w.p. $\geq 1e-4$, Learnt and Entropy PS are in very close agreement. Note that if we were to scramble the message sequence to codeword mapping as proposed in Section III-B to reduce the BLER spread, we must restrict the scrambling to be applied to codewords with the same prior probability, in order to retain the benefits of performing Entropy PS.
@@ IV-C Step PS
While Arithmetic PS enables utilization of a common PS scheme for a range of priors, both Entropy and Arithmetic PS lead to a large variation in power between different codewords. While power shaping was introduced to exploit the non-uniform codeword prior, we would still like to be able to control the power variation. In order to control the power variation, we propose Step PS. Step PS consists of only two power levels – $\alpha_{0}=P_{0}$, power of the all-ACK codeword and $\alpha_{m}=P_{1}$ $\forall m\neq 0$, power of all the other codewords. Furthermore, Step PS can be parameterized by a single value $\delta=P_{1}-P_{0}$. By setting $P_{0}<P_{1}$, Step PS acts as a coarse approximation to Entropy and Arithmetic PS. The power variation in Step PS can be easily reduced by lowering $\delta$, and the signaling overhead is reduced from $2^{k}$ codeword power values to $1$ power differential.
@@
For all PS schemes, the power levels are subjected to the constraint $\sum_{m}\pi_{m}\alpha_{m}=1$. This is to ensure a consistent definition of $\mathrm{SNR}$ during simulation. In practice, such a normalization would not be performed. It would instead be performed implicitly as part of the power control equation in UE behavior (refer Section 7.2.1 in [3gpp.38.213]).
@@ V AWGN Receiver Design
V-A Bitwise UEP
@@
which could be swept to achieve the desired degree of NACK over ACK protection. We will henceforth refer to this decoder as Bitwise UEP. Bitwise MAP is a special case of Bitwise UEP where $\beta/(1-\beta)=p/(1-p)$. From detection theory [steven1993fundamentals], we observe that the likelihood-based criterion in (14) resembles the classical Neyman-Pearson (NP) test. However, the NP test does not require any knowledge of the prior $\mathbf{\Pi}$ since it is applicable for one bit only i.e. $k=1$. In the 1 bit case, (14) simply collapses to comparing $\mathbf{y}$ with a threshold. Depending on the target NACK error rate and the SNR, the threshold can easily be computed (refer Appendix A for details).
Extending the NP test to $k>1$ requires computing the posterior bit probability $p(b_{i}=0|\mathbf{y})$ via Bayes rule as outlined in (11). However, it is unclear how to determine $\beta$ to achieve the desired degree of NACK over ACK protection. We propose an empirical rule to determine the threshold $\beta$ (experimental evidence is presented in Appendix C). Suppose we wish to target an ACK and NACK BER of $\mathrm{BER}_{\mathrm{ACK,t}}$ and $\mathrm{BER}_{\mathrm{NACK,t}}$ respectively. Then, set
@@
V-B MLM GLRT
@@
From (13), it follows that $p_{\Psi_{1}}(\mathbf{y}|b_{i}=j)=\sum_{m\in\mathcal{C}_{b_{i}=j}}p(\mathbf{y}|\mathcal{H}_{m})\pi_{m}(\Psi_{1})/p(b_{i}=j)$.
@@
We refer to this approximation as the Max Log MAP (MLM) approach [hagenauer1994iterative]. MLM GLRT provides two benefits - the ability to compute the bitwise likelihood with partial knowledge of the codeword priors and a reduced complexity implementation (compared to Bitwise UEP). Since we compute an approximation to $p(\mathbf{y}|b_{i}=0)$, the value of $\beta$ as dictated by (15) may no longer be exactly optimal. However, to maintain a fair comparison, we will not change $\beta$ for MLM GLRT when comparing it to Bitwise UEP in Section VIII-F.
@@ VI Fading Channel Receiver Design
In order to compute $p(\mathbf{y}_{l}|b_{i}=1)$, we need to compute $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ accurately.
@@ VI-B Optimal Coherent Receiver
Denote by $\mathbf{\tilde{h}}_{d,l}=\epsilon g\mathbf{h}_{d,l}$. In the equations to follow, we will utilize the subscript $1$ to denote $\mathbf{\tilde{h}}_{d,l}$ and $2$ to denote $\mathbf{y}_{p,l}$ for notational brevity. Conditioned on the $m^{th}$ hypothesis $\mathcal{H}_{m}$, the distribution of $\mathbf{\tilde{h}}_{d,l}$ conditional on $\mathbf{y}_{p,l}=\mathbf{y}_{p}$ is a multivariate Gaussian [eaton2007home] such that
@@
Utilizing the definition of the received signal in (3), we can compute $\mathbf{\Sigma}_{11}$, $\mathbf{\Sigma}_{12}$ and $\mathbf{\Sigma}_{22}$ (refer Appendix D for derivation) as
$\begin{gathered}\mathbf{\Sigma}_{11}=\epsilon^{2}g^{2}\mathbf{R}_{\mathbf{h}_{d}\mathbf{h}_{d}}\\ \mathbf{\Sigma}_{12}=\epsilon^{2}g^{2}\sqrt{\beta_{m}}\mathbf{R}_{\mathbf{h}_{d}\mathbf{h}_{p}}\\ \mathbf{\Sigma}_{22}=\epsilon^{2}g^{2}\beta_{m}\mathbf{R}_{\mathbf{h}_{p}\mathbf{h}_{p}}+\epsilon^{2}\sigma^{2}\mathbf{I}_{N_{p}}.\end{gathered}$ (25)
From (3), we observe that $\mathbf{y}_{d,l}=\sqrt{\alpha}_{m}\mathbf{\tilde{h}}_{d,l}\odot\mathbf{c}_{m}+\epsilon\mathbf{n}_{d,l}$. Hence $\mathbf{y}_{d,l}$ conditioned on $\mathcal{H}_{m}$ and $\mathbf{y}_{p,l}=\mathbf{y}_{p}$ is a complex Gaussian vector with
$\begin{gathered}\boldsymbol{\mu}_{d,l}=\sqrt{\alpha_{m}}\mathbb{E}[\mathbf{\tilde{h}}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l}=\mathbf{y}_{p}]\odot\mathbf{c}_{m}\\ \mathbf{\Sigma}_{d,l}=\alpha_{m}\mathrm{diag}(\mathbf{c}_{m})\mathrm{Cov}[\mathbf{\tilde{h}}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l}=\mathbf{y}_{p}]\mathrm{diag}(\mathbf{c}_{m})^{H}+\epsilon^{2}\sigma^{2}\mathbf{I}_{N_{d}},\end{gathered}$ (26)
@@
From (3), we observe that $\mathbf{y}_{p,l}=\sqrt{\beta_{m}}\mathbf{\tilde{h}}_{p,l}\odot\mathbf{1}+\epsilon\mathbf{n}_{p,l}$. Conditioned on $\mathcal{H}_{m}$, $\mathbf{y}_{p,l}$ is a complex Gaussian vector with
@@
$\arg\max_{m}p(\mathbf{y}_{i}|\mathcal{H}_{m})=\arg\max_{m}p(\mathbf{y}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l}).$ (32)
This explains the rationale for coherent communication in practice – we perform channel estimation and then utilize the channel estimate on the data REs to decode and demodulate the data REs. However, if the DMRS is subject to data dependent power shaping, then $p(\mathbf{y}_{p,l}|\mathcal{H}_{m})$ is no longer independent of $\mathcal{H}_{m}$ and simply computing $p(\mathbf{y}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l})$ is no longer optimal.

### v2

@@ II System Model
Consider a channel code mapping $k$ information bits to $n$ coded symbols, henceforth identified by the tuple $(k,n)$. An information bit value of 0 corresponds to NACK, while 1 corresponds to ACK. A message sequence $\mathbf{b}_{m}=\{b_{0},b_{1},\ldots b_{k-1}\}\in\{0,1\}^{k}$ consisting of $k$ information bits is jointly encoded and modulated to $n$ real values (or equivalently $n/2$ complex values). The message sequence index is denoted by $m$ (for e.g. if $k=4$, then the message sequence 0101 corresponds to index 5). The complex valued encoded sequence is denoted by $\mathbf{c}_{m}$. Each codeword $\mathbf{c}_{m}$ has norm $\|\mathbf{c}_{m}\|^{2}_{2}=n/2$ and is associated with a prior probability $\pi_{m}$ such that $\sum_{m\in[2^{k}]}\pi_{m}=1$. If the encoding is binary, the intermediate encoded bit sequence is denoted by $\mathbf{e}(\mathbf{b}_{m})$. Denote the codebook by $\mathcal{C}=\{\mathbf{c}_{0},\mathbf{c}_{1},\ldots\mathbf{c}_{2^{k}-1}\}$. Subsequently, each codeword $\mathbf{c}_{m}$ is multiplied by a scalar per-codeword power shaping factor $\sqrt{\alpha_{m}}$ (to be introduced in Section IV) such that $\sum_{m}\pi_{m}\alpha_{m}=1$. In current NR specifications [3gpp.38.212], all codewords are transmitted with equal power i.e. $\alpha_{m}=1$.
@@ II-A AWGN setup
where $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{n/2})$ and $\mathrm{SNR}=\sum_{m}\pi_{m}\alpha_{m}g^{2}/\sigma^{2}=g^{2}/\sigma^{2}$.
@@ II-B NR-compliant fading channel setup
Consider the end-to-end (E2E) frequency domain (FD) pipeline for uplink transmission of PUCCH [^3] Note that a similar formulation can be derived for HARQ on PUSCH as well. We focus only on PUCCH for simplicity of the SIMO setup. HARQ-ACK as depicted in Fig. 1. We assume a SIMO setup with $N_{r}$ receive antennas, since PUCCH is transmitted from a single antenna port [johnson20195g]. Under the assumption of coherent reception i.e. excluding PUCCH Format 0, we would also transmit $N_{p}$ Demodulation Reference Symbols (DMRS) resource elements (REs) which we set to 1 w.l.o.g. given a single transmit port. DMRS REs are utilized by the receiver to perform channel estimation (ChEST). Suppose that the DMRS REs are also multiplied by a power shaping factor $\sqrt{\beta_{m}}$. In this paper, we will be considering two cases - no power shaping on the DMRS i.e. $\beta_{m}=1$ and data dependent DMRS power shaping i.e. $\beta_{m}=\alpha_{m}$. In either case, observe that $\sum_{m}\pi_{m}\beta_{m}=1$.
@@
$\mathbf{y}_{l}=g\mathbf{h}_{l}\odot\mathbf{c}_{p+d,m}+\mathbf{n}_{l},$ (3)
where $\mathbf{n}_{l}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I})$ [^4] The system modeling can be extended w.l.o.g. to provide for any receiver scaling that is common to both signal and noise..
Denote the received signal (across all receive antennas) by $\mathbf{Y}=[\mathrm{y}_{1},\ldots\mathrm{y}_{N_{r}}]$. Depending on whether we choose to perform coherent or noncoherent decoding, the received signal is passed to a channel estimator module before being input to a decoder. Furthermore, we can define the long term average SNR (per RE and per receive antenna) and noise variance (NV) measured on the DMRS tones as
$\mathrm{SNR}$ $=\frac{\sum_{m}\pi_{m}\beta_{m}g^{2}}{\sigma^{2}}=\frac{g^{2}}{\sigma^{2}}$ (4)
$\mathrm{NV}$ $=\sigma^{2}.$
@@ III-A Transformer (TF) Architecture
We utilize a transformer (TF) encoder [vaswani2017attention] to encode the input bits. The architecture is illustrated in Fig. 2. The input bit sequence $\mathbf{b}_{m}\in\{0,1\}^{k}$ is divided into $A$ tokens of size $\lceil k/A\rceil$. If $k<A\lceil k/A\rceil$, the first $k$ positions are filled with input bits and the remaining $A\lceil k/A\rceil-k$ positions are set to 0. Subsequently, we BPSK modulate the input bits $\mathbf{m}=(-1)^{\mathbf{b}_{m}}$ before inputting the tokens to the transformer. The transformer encoder consists of $L=6$ layers. Each layer consists of the standard self attention and feedforward networks with residual connections. We utilize an embedding length [^5] Refer https://pytorch.org/docs/stable/generated/torch.nn.TransformerEncoder-Layer.html for meaning of each hyperparameter of $\mathrm{d}_{\mathrm{model}}=128$, feedforward dimension $\mathrm{dim}_{\mathrm{feedforward}}=512$ and $\mathrm{nhead}=8$. The tokens input to the transformer and output by it are passed through Layer Norm [xiong2020layer] normalization. Positional embedding is not added to the input tokens. The transformer outputs $A$ tokens. Each output token is passed through a linear layer $H$ to obtain $\lceil n/A\rceil$ coded symbols per token. The coded symbols computed from the output tokens are concatenated to obtain the codeword of $n$ coded symbols, which can then be normalized to ensure the total codeword power of each codeword is $n/2$. Finally, the transformer output can be reshaped to obtain $\mathbf{c}_{m}$, an $n/2$ length complex codeword.
@@ III-B Codebook Properties
Since the encoding operations involve computing self-attention on the message tokens, the codebook is non-linear. Non-linearity implies that each codeword will have a different Block Error Rate (BLER)-SNR curve, leading to a per-codeword BLER spread.
Since we do not utilize positional embedding, the codebook learnt is group permutation symmetric with a group size of $\lceil k/A\rceil$. This implies we do not need to store the embedding of any group permutations of the input message sequence. Given an input sequence $\mathbf{x}$, we do not need to invoke the transformer to compute the encoding of any sequence that can be written as $\mathbf{Px}$, where $\mathbf{P}\in\{0,1\}^{A\times A}$ and $\mathbf{x}=[[\mathbf{m}_{0}^{T}],[\mathbf{m}_{1}^{T}]\ldots,[\mathbf{m}_{A-1}^{T}]]\in\{-1,1\}^{A\times\lceil k/A\rceil}$. This implies that for any input sequence $\mathbf{x}$, there are at most A!1A!-1 orderings of its input tokens [^6] If some of the tokens are identical, then the number of distinct orderings will be smaller. whose encoding can be computed by group permutation of $H(\mathbf{x})$. This property can be exploited to reduce the size of look-up table (LUT) used to store the codebook. Furthermore, increasing the number of tokens $A$ (or equivalently reducing the token size $\lceil k/A\rceil$) increases the degree of group permutation symmetry in the codebook. In other words, larger symmetry reduces the size of the LUT required to store the codebook (refer Appendix A for further details).
@@ IV-A Entropy PS
Figure 3: Entropy and Learnt PS $\alpha_{m}$ as a function of $\pi_{m}$
Observe that codewords with high probability of occurrence $\pi_{m}$ have low power $\alpha_{m}$ and vice-versa. Moreover, codewords with the same probability of occurrence have the same codeword power. Since we plot as a function of $\pi_{m}$, Entropy PS is a single dot but the Learnt PS has a spread (if more than one codeword has the same prior probability). Since codewords with low $\pi_{m}$ are seen less frequently in training, this is to be expected. However, for all codewords w.p. $\geq 1e-4$, Learnt and Entropy PS are in very close agreement.
@@ IV-C Step PS
While Arithmetic PS enables utilization of a common PS scheme for a range of priors, both Entropy and Arithmetic PS lead to a large variation in power between different codewords. While power shaping was introduced to exploit the non-uniform codeword prior, we would still like to be able to control the power variation. In order to control the power variation, we propose Step PS. Step PS consists of only two power levels – $\alpha_{0}=P_{0}$, power of the all-ACK codeword and $\alpha_{m}=P_{1}$ $\forall m\neq 0$, power of all the other codewords. Furthermore, Step PS can be parameterized by a single value $\delta=P_{1}-P_{0}$, where $P_{0}$ and $P_{1}$ are in dBm and $\delta$ is in dB. By setting $P_{0}<P_{1}$, Step PS acts as a coarse approximation to Entropy and Arithmetic PS. The power variation in Step PS can be easily reduced by lowering $\delta$, and the signaling overhead is reduced from $2^{k}$ codeword power values to $1$ power differential.
@@
For all PS schemes, the power levels are subjected to the constraint $\sum_{m}\pi_{m}\alpha_{m}=1$. This is to ensure a consistent definition of $\mathrm{SNR}$ during simulation. In other words, the long term average SNR in the presence of power shaping In practice, such a normalization would not be performed. It would instead be performed implicitly as part of the power control equation in UE behavior (refer Section 7.2.1 in [3gpp.38.213]).
@@ V AWGN Receiver Design
V-A Bitwise Unequal Error Protection (Bitwise UEP)
@@
which could be swept to achieve the desired degree of NACK over ACK protection. We will henceforth refer to this decoder as Bitwise UEP. Bitwise MAP is a special case of Bitwise UEP where $\beta/(1-\beta)=p/(1-p)$. From detection theory [steven1993fundamentals], we observe that the likelihood-based criterion in (14) resembles the classical Neyman-Pearson (NP) test. However, the NP test does not require any knowledge of the prior $\mathbf{\Pi}$ since it is applicable for one bit only i.e. $k=1$. In the case of repetition coding for $k=1$, (14) simply collapses to comparing $\Re\{\mathbf{1}^{H}\mathbf{y}\}$ with a threshold. Depending on the target NACK error rate and the SNR, the threshold can easily be computed (refer Appendix C).
Extending the NP test to $k>1$ requires computing the posterior bit probability $p(b_{i}=0|\mathbf{y})$ via Bayes rule as outlined in (11). However, it is unclear how to determine $\beta$ to achieve the desired degree of NACK over ACK protection. We propose an empirical rule to determine the threshold $\beta$ (experimental evidence is presented in Appendix D). Suppose we wish to target an ACK and NACK BER of $\mathrm{BER}_{\mathrm{ACK,t}}$ and $\mathrm{BER}_{\mathrm{NACK,t}}$ respectively. Then, set
@@
V-B Max Log MAP Generalized Likelihood Ratio Test (MLM GLRT)
@@
From (13), it follows that $p_{\Psi_{1}}(\mathbf{y}|b_{i}=j)=\sum_{m\in\mathcal{C}_{b_{i}=j}}p(\mathbf{y}|\mathcal{H}_{m})\pi_{m}(\Psi_{1})/p(b_{i}=j)$ [^7] Observe that this does not capture the dependence of the marginal probability $p(b_{i}=j)$ on $\Psi_{1}$. This is valid only if $p(b_{i}=j)$ is independent of $\Psi_{1}$ as will be considered in Section VII-A..
@@
We refer to this approximation as the Max Log MAP (MLM) approach [hagenauer1994iterative]. MLM GLRT provides two benefits - the ability to compute the bitwise likelihood with partial knowledge of the codeword priors and a reduced complexity implementation (compared to Bitwise UEP). Since we compute an approximation to $p(\mathbf{y}|b_{i}=0)$, the value of $\beta$ as dictated by (15) may no longer be exactly optimal i.e. achieve the target ACK and NACK error rates at the same SNR. However, in the absence of (15), an optimal value for $\beta$ cannot be determined in practice since the gNB is unable to ascertain if the received signal has been correctly decoded owing to a lack of CRC in the short block-length regime. Hence, to maintain a fair comparison, we will not change $\beta$ for MLM GLRT when comparing it to Bitwise UEP in Section VIII-F.
@@ VI Fading Channel Receiver Design
In order to compute $p(\mathbf{y}_{l}|b_{i}=1)$, we need to compute $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ accurately. While doing so, we can either perform joint channel estimation and detection, which we will refer to as a noncoherent receiver design (since we do not explicitly estimate the channel) or we can first estimate the channel from the pilot REs and then perform detection as part of a coherent receiver design.
@@ VI-B Optimal Coherent Receiver
Denote by $\mathbf{\tilde{h}}_{d,l}=g\mathbf{h}_{d,l}$. In the equations to follow, we will utilize the subscript $1$ to denote $\mathbf{\tilde{h}}_{d,l}$ and $2$ to denote $\mathbf{y}_{p,l}$ for notational brevity. Conditioned on the $m^{th}$ hypothesis $\mathcal{H}_{m}$, the distribution of $\mathbf{\tilde{h}}_{d,l}$ conditional on $\mathbf{y}_{p,l}=\mathbf{y}_{p}$ is a multivariate Gaussian [eaton2007home] such that
@@
Utilizing the definition of the received signal in (3), we can compute $\mathbf{\Sigma}_{11}$, $\mathbf{\Sigma}_{12}$ and $\mathbf{\Sigma}_{22}$ (refer Appendix E for derivation) as
$\begin{gathered}\mathbf{\Sigma}_{11}=g^{2}\mathbf{R}_{\mathbf{h}_{d}\mathbf{h}_{d}}\\ \mathbf{\Sigma}_{12}=g^{2}\sqrt{\beta_{m}}\mathbf{R}_{\mathbf{h}_{d}\mathbf{h}_{p}}\\ \mathbf{\Sigma}_{22}=g^{2}\beta_{m}\mathbf{R}_{\mathbf{h}_{p}\mathbf{h}_{p}}+\sigma^{2}\mathbf{I}_{N_{p}}.\end{gathered}$ (25)
From (3), we observe that $\mathbf{y}_{d,l}=\sqrt{\alpha}_{m}\mathbf{\tilde{h}}_{d,l}\odot\mathbf{c}_{m}+\mathbf{n}_{d,l}$. Hence $\mathbf{y}_{d,l}$ conditioned on $\mathcal{H}_{m}$ and $\mathbf{y}_{p,l}=\mathbf{y}_{p}$ is a complex Gaussian vector with
$\begin{gathered}\boldsymbol{\mu}_{d,l}=\sqrt{\alpha_{m}}\mathbb{E}[\mathbf{\tilde{h}}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l}=\mathbf{y}_{p}]\odot\mathbf{c}_{m}\\ \mathbf{\Sigma}_{d,l}=\alpha_{m}\mathrm{diag}(\mathbf{c}_{m})\mathrm{Cov}[\mathbf{\tilde{h}}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l}=\mathbf{y}_{p}]\mathrm{diag}(\mathbf{c}_{m})^{H}+\sigma^{2}\mathbf{I}_{N_{d}},\end{gathered}$ (26)
@@
From (3), we observe that $\mathbf{y}_{p,l}=\sqrt{\beta_{m}}\mathbf{\tilde{h}}_{p,l}\odot\mathbf{1}+\mathbf{n}_{p,l}$. Conditioned on $\mathcal{H}_{m}$, $\mathbf{y}_{p,l}$ is a complex Gaussian vector with
@@
$\arg\max_{m}p(\mathbf{y}_{l}|\mathcal{H}_{m})=\arg\max_{m}p(\mathbf{y}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l}).$ (32)
Note that $p(\mathbf{y}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l})$ is computed by first computing the LMMSE channel estimate $\hat{\mathbf{h}}_{l}=\sqrt{\alpha_{m}/\beta_{m}}\mathbf{W}_{m}\mathbf{y}_{p}$ and then substituting it in (29). This motivates the rationale behind coherent communication in practice – we perform channel estimation and then utilize the channel estimate on the data REs to decode and demodulate the data REs [^8] Note that (29) computes $\Sigma_{d,l}$ optimally as well which might not be done in practice.. However, if the DMRS is subject to data dependent power shaping, then $p(\mathbf{y}_{p,l}|\mathcal{H}_{m})$ is no longer independent of $\mathcal{H}_{m}$ and simply computing $p(\mathbf{y}_{d,l}|\mathcal{H}_{m},\mathbf{y}_{p,l})$ is no longer optimal.